OpenStudy (anonymous):

write an equation of the line containing the given point and parellel to the give line. express your answer in the form y=mx+b (-9,9);6x=5y+17

OpenStudy (amistre64):

parallel lines share the same x and y baggage; perps switch out the coeefs

OpenStudy (amistre64):

so get rid of that +17 and fill in with your new point to get a replacement for it

OpenStudy (amistre64):

(-9,9) 6x=5y+17 6x=5y+n 6(-9)=5(9)+n -54 =45 +n -99 = n

OpenStudy (amistre64):

6x+99 ------ = y 5